OpenStudy (anonymous):

Will give Medal! At the beginning of the week, Paloma opened a new box of 125 coffee creamers. Sixty-one coffee creamers were used during the week. Which equation could be used to find x, the number of coffee creamers left at the end of the week? 61 - 125 = x x - 61 = 125 x + 61 = 125 61x = 125

OpenStudy (anonymous):

@Michele_Laino @mathmath333 @iambatman @inkyvoyd @kittiwitti1

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@LittleRedRiding

OpenStudy (anonymous):

Can you help me?

OpenStudy (littleredriding):

umm...can you write the answers out more clearly?

OpenStudy (igreen):

Which one tells you to subtract 125 - 61?

OpenStudy (anonymous):

Oops it logged me out so for get this question and I will have a diffrent one.

OpenStudy (anonymous):

Amy and Stephanie were playing a number game. Amy picked a number and told Stephanie that if she added 8 to the number and multiplied the sum by 3, it would equal 9. Which equation could Stephanie use to find the number, x? 3 + (x + 8) = 9 3(x) + 8 = 9 8(x + 3) = 9 3(x + 8) = 9

OpenStudy (igreen):

First you have to add 8 to the number..then you multiply 3..which one tells you to do that?

OpenStudy (kittiwitti1):

Please don't bribe for answers. Also, please post your questions in a new post each, so we have an easier time helping you answer them. (: So, we have x. Add 8 to x: 8+x Then multiply the sum by 3. The sum would be (8+x), so 3(8+x). Answer: 3(x+8)=9

OpenStudy (anonymous):

So the 4th one

OpenStudy (igreen):

@kittiwitti1 Please don't give answers. We are here to guide people to the answer, so that they get it on their own.

OpenStudy (kittiwitti1):

*I mean, what do you think? ^^;

OpenStudy (anonymous):

Thank you! I have another one.

OpenStudy (kittiwitti1):

I did explain the process @iGreen... I know better than to regurgitate answers...

OpenStudy (anonymous):

Samantha donated 9 T-shirts to charity and has 15 left. Which equation could be used to find x, the number of T-shirts Samantha had originally? x - 15 = 9 x - 9 = -15 x + 9 = 15 x + 15 = -9

OpenStudy (kittiwitti1):

Please post your question in a new post @HappySprout (:

OpenStudy (igreen):

@kittiwitti1 Yes, but someone can easily just take your answer and not read your explanation. Give the explanation and let the asker get the answer on his/her own.

OpenStudy (anonymous):

Oh sorry ok I will do that now

OpenStudy (kittiwitti1):

Alright @iGreen Thank you @HappySprout :)

OpenStudy (anonymous):

Thank you!

OpenStudy (kittiwitti1):

You're very welcome @HappySprout (: We're here to help! ^^